Ball A travelling with velocity v collides head on with a stationary ball B. What is the velocity of ball B immediately after impact if the two balls have equal masses?


1
Expert's answer
2021-05-23T16:33:56-0400

From the law of conservation of momentum, we have:


"m_Av_{A,i}=m_Av_{A,f}+m_Bv_{B,f}."

Let's suppose that the kinetic energy is conserved. Then, we can write:


"\dfrac{1}{2}m_Av_{A,i}^2=\dfrac{1}{2}m_Av_{A,f}^2+\dfrac{1}{2}m_Bv_{B,f}^2."

This formula gives us an addititional relationship between velocities. Therefore, with the help of these two formulas we can find the velocity of ball B immediately after impact. Since "m_A=m_B=m" and "v_{A,i}=v", we get:


"v_{B,f}=\dfrac{2m_Av_{A,i}}{(m_A+m_B)}=\dfrac{2mv}{2m}=v."